if sin a = 12 /13 ,cos b = 3/5 AND NEITHER a nor b is in first quadrant ,then find quadrant in which a+b lies

Grade:11

2 Answers

Arun
5 years ago
Dear student
 
sin a = 12/13
 
Hence a will be in 2nd quadrant.
 
cos b = 3/5 
Hence b will lie in 4th quadrant
 
Now
 
 sin (a+b) = sin a cos b + cos a sin b
 
 = (12/13) * (3/5) + (-5/13) *( -4/5)
 
 = 56/65
 
 cos (a +b) = -15/65 + 48/65 = 33/65
 
Hence (a+b ) will lie in 1st quadrant
Sourab
5 years ago
Yo yo bantai rapper 
Charas Ganja mere Ko pyara
Insan hi Insan Ko Manta nahi
Aye Rupali aye Rupali 
Carry Minati is the best one watch on YouTube broooooo
 
 
A+B will lie on 1st quadrant